# Scalable analysis of stop-and-go waves

### v1

@@ 8 Conclusion
This article presents an automatic and scalable approach for identifying the stop-and-go wave boundaries, enabling the analysis of stop-and-go waves at scale. The proposed method is capable of capturing the wave generation, propagation, dissipation, as well as bifurcation and merging. It is built upon a graph-based representation of the spatio-temporal points associated with stop-and-go waves, specifically wave front (start) points and wave tail (end) points, and approaches the solution as a graph component identification problem. We implement it in Python and demonstrate it on a large-scale dataset, I-24 MOTION INCEPTION. New insights revealed from this demonstration with emerging phenomena include: (a) we demonstrate that waves do generate, propagate, and dissipate at a scale (miles and hours) and ubiquity never observed before; (b) wave fronts and tails travels at a consistent speed for a critical speed between 10-20 mph, with propagation variation across lanes, where wave speed on the outer lane are less consistent compared to those on the inner lane; (c) wave fronts and tails propagate at different speeds; (d) wave boundaries capture rich and non-trivial wave topologies (with several merges and bifurcations per wave boundary on average), highlighting the complexity of waves.
The scale and the complexity of the stop-and-go waves may encourage researchers in the community to revisit the stop-and-go waves phenomena observed in the NGSIM dataset and other experimental datasets. The rich topologies within the wave boundaries warrant further investigation.
@@
The code developed in this article will be available at https://i24motion.org. The tools introduced in this article could open new possibilities for analyzing stop-and-go traffic waves at scale and it could lower the barrier for researchers to deal with large-scale trajectory data. For example, the I-24 MOTION instrument [9] will enable experiments involving variable speed limit control and lane-control systems. The method and tool make it feasible to analyze the impact of the control strategies on the traffic wave dynamics.
Acknowledgement
@@
The authors would like to extend their gratitude to Professor Martin Treiber at TU Dresden for his insights regarding the dimensions of Edie’s box. This work is supported by the Tennessee Department of Transportation (TDOT) under Grant No. OTH2023-01F-01 and the National Science Foundation (NSF) under Grant No. 2135579 (Work, Sprinkle), and 2111688 (Sprinkle). The views expressed herein do not necessarily represent the views of the Tennessee Department of Transportation or the the United States Government.
@@ A.1 Virtual sensors and virtual vehicles
Long-range raw vehicle trajectory data is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [97; 98]. The tasks fail when errors occurring at the current frame cannot be corrected using information from subsequent frames. These errors may arise from various system sources, including homography estimation [97] and object occlusion [99]. However, these challenges do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. In this context, generating vehicle trajectories [105] from high-fidelity speed field can serve as a surrogate measure for the raw trajectory data, enabling the analysis of stop-and-go waves.We use the virtual trajectory generation method from [100] as a first step in wave identification in this work. We detail this approach below. Table 3 provides a comprehensive list of the variables and parameters used throughout the appendix.
@@
$N$ total number of trajectories from raw data - -
@@
$N$ the number of virtual vehicles - -
@@ A.1.1 Virtual sensors: Edie’s definition for macroscopic measurements
Edie [106] provides an approach to calculate spatio-temporal mean of density, flow speed, and traffic flow from vehicle trajectories, which is named as Edie’s definition (shown in Figure 11). According to the definition, the density and the flow can be computed from the total travel time (TTT) and total travel distance (TTD) of all the vehicles within an area. Using Edie’s definition allows for the parallel calculation of measurements on a trajectory-by-trajectory basis [100]. Consider a shear box [107; 105] centered at a point ($t$, $x$), and let $\Delta t$ and $\Delta x$ denote the height and width of the box, as shown in Figure 11. The macroscopic estimates can be computed as:
@@ A.1.2 Adaptive Smoothing
With the original speed field generated, virtual trajectories can be created. However, the original macroscopic measurements often contain missing values and outliers. As suggested by [105], an adaptive smoothing method should be applied to the raw data to improve accuracy. The Adaptive Smoothing Method (ASM), developed by [108; 109], is a widely used smoothing and interpolation algorithm for constructing a continuous spatio-temporal mean speed field. While it is especially useful for data collected from fixed infrastructure sensors, such as inductive loops, it is also applicable to the macroscopic speed data generated from trajectories. This method is particularly helpful in interpolating speeds in occluded areas, such as under bridges. We directly apply this method, with parameter settings as listed in Table 9.
@@
(a) Congestion kernels
(b) Free-flow kernels
Figure 12: Weights of homogeneous filters: Figure 12(a) with congestion and Figure 12(b) with free-flow kernels. In a matrix setting, this is consistent with the Figure 1 in [108].
Since the primary goal of applying ASM in this paper is interpolation, a small smoothing kernel is selected, with the kernel weights illustrated in Figure 12. The code and numerical kernel weights used are available at https://i24motion.org.
@@ 1 Initialize a list of components $\mathcal{C}^{d}$ and $i=0$;
Initialize a set $\mathcal{V}_{v}\leftarrow\emptyset$ To keep track of visited nodes
@@ 29            Add $\mathcal{C}_{k}$ to $\mathcal{C}$;
Appendix C Wave fronts and tails travel distance and speed

### v2

@@ 8 Conclusion
This article presents an automatic and scalable approach for identifying stop-and-go wave boundaries, enabling large-scale analysis of such phenomena. The method captures wave generation, propagation, dissipation, as well as bifurcation and merging. It builds on a graph-based representation of the spatio-temporal points associated with stop-and-go waves, specifically wave front (start) points and wave tail (end) points, and frames the solution as a graph component identification problem. This study establishes a foundation for measuring traffic wave properties, a prerequisite for evaluating how microscopic [116], mesoscopic [117], and macroscopic [118] models reproduce observed traffic wave phenomena. The approach can also serve as a performance metric for iterative calibration and hyper-parameter optimization. Recent applications, including the use of generative artificial intelligence for traffic wave reconstruction [119], have applied this method to measure the physical properties of generative outputs.
In addition, this study provides new empirical insights into the complexity of traffic wave dynamics. By identifying and clustering approximately 400 waves per lane through a graph-based representation, we demonstrate for the first time that traffic waves can bifurcate and merge at scales not previously observed. These findings highlight the need for traffic models to incorporate bifurcation and merging phenomena in order to more realistically capture traffic flow dynamics. While road geometry plays a role in these dynamics, our analysis shows it is not the sole determining factor, suggesting a more combined effects of different factors. The scale and the complexity of the stop-and-go waves may encourage researchers in the community to revisit the stop-and-go waves phenomena observed in the NGSIM dataset and other experimental datasets. The rich topologies within the wave boundaries warrant further investigation.
@@
The code developed in this article is publicly available at https://github.com/I24-MOTION/wave-analysis and will be listed on https://i24motion.org. The tools introduced in this article could open new possibilities for analyzing stop-and-go traffic waves at scale and it could lower the barrier for researchers to deal with large-scale trajectory data. For example, the I-24 MOTION instrument [11] will enable both CAV field experiments [81; 120] and ITS deployments [121]. It is worthy of attention that the method and tool presented in this article allow for empirically analyzing how control strategies and vehicle automation influence traffic wave properties, a growing topic within traffic flow theory and characteristics [122; 123; 124; 93; 119; 120]. Further research will focus on investigating the factors influencing wave evolution [65], accompanied by refinements in data quality.
Acknowledgment
@@
The authors would like to extend their gratitude to Professor Martin Treiber at TU Dresden for his insights regarding the dimensions of Edie’s box. Preliminary version of this work is presented at Conference in Emerging Technologies in Transportation Systems (TRC-30). This work is supported by the Tennessee Department of Transportation (TDOT) under Grant No. OTH2023-01F-01 and the National Science Foundation (NSF) under Grant No. 2135579 (Work, Sprinkle), and 2111688 (Sprinkle). The views expressed herein do not necessarily represent the views of the Tennessee Department of Transportation or the United States Government.
@@ A.1 Virtual sensors and virtual vehicles
Long-range raw vehicle trajectory data is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [103; 104]. The tasks fail when errors occurring at the current frame cannot be corrected using information from subsequent frames. These errors may arise from various system sources, including homography estimation [103] and object occlusion [105]. However, these challenges do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. In this context, generating vehicle trajectories [125] from high-fidelity speed field can serve as a surrogate measure for the raw trajectory data, enabling the analysis of stop-and-go waves.We use the virtual trajectory generation method from [106] as a first step in wave identification in this work. We detail this approach below. Table 3 provides a comprehensive list of the variables and parameters used throughout the appendix.
@@
NN total number of trajectories from raw data - -
@@
NN the number of virtual vehicles - -
@@ A.1.1 Virtual sensors: Edie’s definition for macroscopic measurements
Edie [126] provides an approach to calculate spatio-temporal mean of density, flow speed, and traffic flow from vehicle trajectories, which is named as Edie’s definition (shown in Figure 15). According to the definition, the density and the flow can be computed from the total travel time (TTT) and total travel distance (TTD) of all the vehicles within an area. Using Edie’s definition allows for the parallel calculation of measurements on a trajectory-by-trajectory basis [106]. Consider a shear box [127; 125] centered at a point ($t$, $x$), and let $\Delta t$ and $\Delta x$ denote the height and width of the box, as shown in Figure 15. The macroscopic estimates can be computed as:
@@ A.1.2 Adaptive Smoothing
With the original speed field generated, virtual trajectories can be created. However, the original macroscopic measurements often contain missing values and outliers. As suggested by [125], an adaptive smoothing method should be applied to the raw data to improve accuracy. The Adaptive Smoothing Method (ASM), developed by [128; 129], is a widely used smoothing and interpolation algorithm for constructing a continuous spatio-temporal mean speed field. While it is especially useful for data collected from fixed infrastructure sensors, such as inductive loops, it is also applicable to the macroscopic speed data generated from trajectories. This method is particularly helpful in interpolating speeds in occluded areas, such as under bridges. We directly apply this method, with parameter settings as listed in Table 9.
@@ 1 Initialize a list of components $\mathcal{C}^{d}$ and $i=0$;
2
Initialize a set $\mathcal{V}_{v}\leftarrow\emptyset$ # To keep track of visited nodes
@@ 29    Add $\mathcal{C}_{k}$ to $\mathcal{C}$;
Appendix C The size of spatio-temporal search box
Before diving into the details, the key takeaway is that if the box size is too small, the method fails, whereas a larger box size remains effective. We recommend setting the temporal size to 3 times the sending frequency of virtual vehicles, with the spatial size adjusted accordingly. Our method is computationally efficient and not highly sensitive to box size. Each run for a single day and lane is completed within 60 seconds when tested on a 2022 MacBook Air with an Apple M2 chip and 16GB of memory.
First, we aim to provide details on the choice of the box size in our paper. In this study, the spatio-temporal box is defined as 20 seconds by 0.07 miles, extending spatially from -0.02 miles to 0.05 miles and temporally from -5 seconds to 15 seconds, as illustrated in Figure 16(a). The - sign indicates upstream in space and the past in time. The look-ahead box size is determined based on wave travel speed and expected time headway, ensuring it effectively captures the Newell-like car-following characteristics of the next vehicle.
Figure 16: Sensitivity analysis of the spatio-temporal box: The effectiveness of our method deteriorates when the box size is too small.
To achieve this, we use 3 times the sending frequency of virtual vehicles. The backward extension in time and space accounts for scenarios where waves remain stationary [112] or when drivers engage in preventive driving, causing braking points to occur even earlier than those of the preceding vehicle. In this paper, virtual vehicles are sent to the mean speed field at 5-second intervals to generate virtual trajectories. A 15-second look-ahead window is then applied to identify the next vehicle. Starting from vehicle index $i$, we check vehicle $i+1$ only. If there are any points from other vehicle indices, such as $i+k$ where $k\geq 2$, we do not attempt to find them. If there are multiple nodes found in the spatio-temporal box, it will connect to the one closest in time. A toy example is shown in Figure 16(b), where node $a$ and $b$ are located within the box. The edge will be connected to $a$, as it is closer in time.
Appendix D Comparison with other traffic wave identification methods
In the literature, identifying traffic waves typically involves two steps: (a) detection (e.g., frequency domain method) and (b) grouping (e.g., clustering) [31; 93]. The detection method we use is based on the critical speed approach [4; 34] and the grouping method we propose involves a graph representation and its connected component identification. Our proposed method contributes the graph topological analysis of traffic waves at the grouping step, offering a novel perspective. In comparison, existing methods are limited in their ability to describe wave bifurcation and merging, as summarized in Table 2.
  1. (a)
    Detection: Wavelet transform and critical speed method.
    We would like to clarify that we do not claim our detection method is novel or better than the others. We like the critical speed method due to its straightforward definition, which corresponds to the speed contour line in the space-time diagram [4].
    A different approach is the Wavelet transform [20], which is a frequency-domain technique for detecting stop-and-go waves. We compare the Wavelet transform with the critical speed method to explain the reasons for selecting the critical speed method in Figure 18. There are several parameters and hyper-parameters in the Wavelet transform, including the mother wavelet, scale parameters (minimum and maximum of scale), and the width of spread. We use the pywt Python package [131] to implement the code, and the parameter we choose it the gaus2 mother wavelet, min_scale=1, max_scale=32, width=2.
    Figure 17: The second-order Gaussian mother wavelet.
    The gaus2 wavelet is visualized in Figure 17, with a pattern of slow down then speed up in amplitude. We then determine the maximum coefficient for each scaled signal to identify the best-fitting frequency and distinguish the front and tail based on the sign (+/-) of the coefficient. The detected wave fronts/tails are shown in Figure 18.
    Figure 18: Demonstration of Wavelet transform in detecting the wave fronts and tails: The top diagram represents the speed-time graph with 10 sampled trajectories, where the wave front is marked in red and the tail in green. The bottom diagram illustrates the space-time diagram with all the key spatiotemporal features labeled as well. The data here is from 2022-11-22, covering virtual vehicle indices 380 to 390.
    The Wavelet transform offers the advantage of enabling multi-scale signal analysis, but it requires careful selection of the mother wavelet and fine-tuning of its parameters. As can be seen from Figure 18, a 10- vehicle platoon experiences multiple traffic oscillations, and each oscillation has different patterns, because the traffic waves are at different stages of development. Tuning the wavelet parameters individually for each wave becomes challenging when applying the technique to large datasets with many types of waves. We thus opt for the critical speed method used in our work to simplify the detection step, as it requires only a single parameter to control the performance.
  2. (b)
    Grouping: DBSCAN and our proposed graph representation.
    DBSCAN is a widely used clustering method [31; 93] for grouping spatio-temporal points in traffic wave analysis. While it performs well, its limitation lies in its inability to identify the topology of the waves. We replicate the methods in [93] and use the sklearn Python package to implement.
    We make slight adjustments in the definition to waves to ensure a fair comparison with our method. In [93], waves are defined as instances where the speed has decreased by at least 10% compared to five seconds earlier. In our case, we define this as the speed falling below the critical speed, using 15 mph as an example.
In [93], the distance for clustering is redefined as below.
$d((t_{1},x_{1},y_{1}),(t_{2},x_{2},y_{2}))=\begin{cases}0,&|x_{1}-x_{2}|\leq x_{d}\land|t_{1}-t_{2}|\leq t_{d}\land y_{1}=y_{2}\\ 1,&\text{Otherwise}\end{cases}$ (21)
where the $t$, $x$, $y$ denote time, space, and lane, respectively. As can be seen from the definition, it closely resembles the spatio-temporal search box described in our paper. If the spatial difference $|x_{1}-x_{2}|$ is within a threshold $x_{d}$, the temporal difference $|t_{1}-t_{2}|$ is within a threshold $t_{d}$, and both points are in the same lane $y_{1}=y_{2}$, the function returns 0, indicating closeness. Otherwise, it returns 1, indicating separation. For a fair comparison, we choose a box size similar to that used in our method, with $x_{d}=0.05$ miles and $t_{d}=15$ seconds.
(a) (a) Our method
(b) (b) DBSCAN
Figure 19: Comparison of our method and the DBSCAN clustering method: (a) Our method identifies a set of wave fronts and tails, along with their connections, allowing us to construct wave skeleton lines (blue) for a structured representation. For detailed visualization, click on https://trafficwaves.github.io/2022-11-22/lane_1/4_15.svg. (b) DBSCAN identifies a set of points under the critical speed.
In Figure 19, we selected the same wave identified by both our method and DBSCAN. While DBSCAN provides a set of points within the wave, our method is specifically designed to define the wave boundary and its topology. Our method automatically identifies how many wave fronts and tails are in this wave component, wave bifurcation and merge details, specifying which waves it bifurcates into, as well as the time and location of bifurcation and merging points. While manual inspection from DBSCAN can also determine the time and space of bifurcation and merging from DBSCAN, our method automates this process, making it a scalable solution for analyzing traffic waves from multiple days of data. In summary, the main benefit of our method is to capture these topological elements automatically, which is not done in clustering approaches such as DBSCAN.
Appendix E Wave fronts and tails travel distance and speed